OpenStudy (anonymous):

A towns high temperature for a given month is 88F and the low temperature is 65F. Write an inequality to represent the range of temperatures for the given month.

OpenStudy (anonymous):

this highest it can be is equal to or less than 88 so \[x \le 88\] and it can be greater than or equal to 65 \[65 \le x\] put them together and it looks like what?

OpenStudy (anonymous):

uhh.. idk.

OpenStudy (anonymous):

i'm not to give you answers so see if this helps 65 = D 88 = W \[D \le x \le W\] can you do it now?

OpenStudy (anonymous):

um ok? and yeah. 65<= x <= 88.

OpenStudy (anonymous):

yeah